solve the logarithymatic equation. Express irratiional solitions in simplist radical form.
log(x+1)-log(x+5)=log1/x

Answer by stanbon(29603)

solve the logarithymatic equation. Express irratiional solitions in simplist radical form.
log(x+1)-log(x+5)=log(1/x)
log(x+1) = log(1/x)+log(x+5)
log(x+1) = log[(1/x)(x+5)]
Take the anti-log to get:
(1/x)(x+5) = x+1
1 + (5/x) = x+1
Multiply thru by x to get:
x + 5 = x^2 + x
x^2=5
x = sqrt(5) or x = -sqrt(5)
------------------------------
-sqrt(5) is extraneous as it would make log(x+1) be the log of a negative #.
------------------------------
Ans: x = sqrt(5)
==================================
Cheers,
Stan H.